how to solve y'''+2y''+y'=10


1
Expert's answer
2021-06-17T14:49:07-0400

Solution.

y"'+2y''+y'=10

"\lambda^3+2\lambda^2+\lambda=0"

"\lambda(\lambda^2+2\lambda+1)=0"

"\lambda_1=0,\lambda_2=-1,\lambda_3=-1."

From here solution of homogeneous equation "y'''+2y''+y'=0" is

"y=C_1+C_2e^{-t}+C_3te^{-t},"

where "C_1,C_2,C_3" are some constants.

Find particular solution of equation "y'''+2y''+y'=10" in the form "y=At+B."

Then

"y'=A, y''=y'''=0."

From here "A=10,B=0."

We will have solution

"y=C_1+C_2e^{-t}+C_3te^{-t}+10t."

Answer. "y=C_1+C_2e^{-t}+C_3te^{-t}+10t."